OpenStudy (anonymous):

AWARDING MEDALS! Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences. F(x) = (x-1) (x-2)

OpenStudy (anonymous):

yes it is a good method....lol

OpenStudy (anonymous):

If the discriminant is negative, the equation has no solution. Thus, completing the square would not help you.

OpenStudy (anonymous):

http://www.khanacademy.org/math/trigonometry/polynomial_and_rational/quad_formula_tutorial/v/solving-quadratic-equations-by-square-roots What does f(x) mean? 4 Votes • 2 Comments • Flag 6 months ago by Umer Naeem f(x) is just the fancy term for y as Sal has mentioned in some of his videos f(x) or y is a function where in if we plug in a value of x it would map to a coordinate of y. for instance if we have a function: f(x)=x+7 If we plug in 3 for x f(3)=3+7 f(3)=10 So our coordinates (x,y) are (3,10) 6 Votes • Comment • Flag 6 months ago by Petrie (Peter S. Asiain III) Show all 3 answers • Answer this question Where did you get the 8 in the beginning of the video where you make each side an 8 4 Votes • Comment • Flag 8 months ago by Cassidy Shuler He is adding 8 to each side of the equation to 'solve' for the (x+1)^2 first. 8 Votes • Comment • Flag 8 months ago by Bollen.DCS Show all 2 answers • Answer this question so you couldnt write (x^2 +81) for (x+9)^2? im stupid 2 Votes • Comment • Flag 6 months ago by maxium213 It is a very common mistake among beginners to write: (x+9)^2 = x^2 + 81. Most of the time people realize this mistake if they get into the habit of writing: (x+9)^2=(x+9)(x+9). Then multiplying this using the distributive property (FOIL), you get: (x+9)^2 = x^2 + 9x + 9x + 81 = x^2 + 18x + 81. 8 Votes • Comment • Flag 6 months ago by MrJones Show all 2 answers • Answer this question In problem number one, how does Sal know what x is to equal zero? 3 Votes • Comment • Flag 2 years ago by Samantha Munsell Because a quadratic equation always equals zero. If you take the factored form of a quadratic, then they will both equal zero, rearrange to make x= and you get both the solutions to the equations. Quadratic equations always equalling zero is a fundamental of quadratics. :) 5 Votes • 1 Comment • Flag 8 months ago by Will Barnard Answer this question... How would you solve -8n(10n-1)=0 using the zero-product property? I'm not really getting exactly how to use this process. Could someone help me with this? 2 Votes • Comment • Flag 7 months ago by Emilia The point of the zero-product property is this: If two or more factors are multiplied together to make 0, then one of the factors must = 0. Think about it, if you want to make 0 by multiplying, you have to have a 0 as a factor: 0 * 8 = 0 125 * 0 = 0 1/4 * 0 = 0 2 * 3 * 4 * 5 * 6 * 0 = 0 So, if you know that a product is 0, then one of the factors MUST be 0. In this problem: -8n(10n-1)=0 The product is 0. The two factors are: -8n and 10n-1 At least one of the factors must be =0. So, set each one equal to 0. -8… (more) 6 Votes • Comment • Flag 7 months ago by Sandy Knight Answer this question... how would u solve this: (x+2)^2=10 2 Votes • 3 Comments • Flag 6 months ago by leann lawer There are two possible answers here, namely: Answer using the positive + square root:x=-2+√10 Answer using the negative - square root:x=-2-√10 Solution: (x+2)²=10 → this is our original equation √(x+2)²=√10 → let's take the square roots of both sides x+2=±√10 → this is what we'll get, then I'll subtract 2 from both sides taking the positive square root..... x=√10-2 → this is what I got, and now I'll rearrange the order x=-2+√10 → I rearranged the order and there we have it! or taking the negative… (more) 6 Votes • 6 Comments • Flag 6 months ago by Petrie (Peter S. Asiain III) Answer this question... I have a problem. I have home work with the same name on this video. But i still dont get this problem... \ x+e-26 = m squared i need answer's 5 Votes • Comment • Flag 8 months ago by Shania Greenwell Does 'e' in your homework problem stand for the mathematical constant? 1 Vote • Comment • Flag 6 months ago by Gabriel Savage Show all 2 answers • Answer this question how did he know to add(+) or subtract( -) 1 on both sides? What's the rule? Thx 3 Votes • Comment • Flag 2 months ago by Joy Mo The rule is that if you add something to one side you have to add it to the other side too. If you subtract something from one side, you have to subtract it from the other side. What you choose will depend on what you want to do. For instance, if there is a -5 on the left side and you want it on the right side, you add 5 to both sides. 3 Votes • Comment • Flag 2 months ago by Sid Answer this question... 3:11 sal puts -1 in front of the plus or minus 2 sq 2. up to now i remember that to shift to the other side of the equation you must acting upon both sides. ex: plus or minus 2 sq 2 MINUS 1. what does he mean when he says one cannot figure this out without a calc? what law allows sal to plus or minus 2 sq 2 from -1? from what i remember, there is no associative law of addition. 1 Vote • Comment • Flag 3 days ago by powerfulwizard99 +2√2 = 2.83 -2√2 = -2.83 +2√2 - 1 = 2.83 - 1 = 1.83 = -1 + 2.83 = -1 + 2√2 -2√2 - 1 = -2.83 - 1 = -3.83 = -1 - 2.83 = -1 - 2√2 3 Votes • Comment • Flag 3 days ago by Sid Show all 2 answers • Answer this question